OpenStudy (anonymous):

49(r-4)^2+70(r-4)+25 = ?????????

OpenStudy (anonymous):

\[49 (r-4)^2+70 (r-4)+25=(7 r-23)^2 \]

OpenStudy (anonymous):

No .. look at this .. let \[\left( r-2 \right) = x\] then we have : \[49x ^{2}+70x+25\] \[=\left( 7x+5 \right)^{2}\] \[=\left[ 7\left( r-2 \right)+5 \right]^{2}\]

OpenStudy (anonymous):

@ Abdullah Al-Hagbani Your "let" expression is incorrect for this problem. It should have been: let (r-4) = 2